# Vector Quantized-Elites: Unsupervised and Problem-Agnostic Quality-Diversity Optimization

### v1

@@ IV-B Evolution Phase
The next key component of VQ-Elites is the evolution phase. During this phase the Quality-Diversity algorithm is executed, performing the usual steps. Foremost, the archive is randomly initialized with random solutions. Then, a random batch of solutions is selected, which are alternated through mutation and crossover. The new variation of the solutions are executed, and evaluated through their behavioral descriptor. In order to extract these BDs, we encode the raw data produced by our solutions with the encoder of our VQ-VAE model, acquiring the corresponding continuous latent vectors. For each BD, we find the closest grid cell, and if the cell is empty the solution is added to the archive, otherwise, we compare the fitness of the current solution with the one of the solution that is already in the archive. The solution in the archive is replaced only if the new solution is fitter. This procedure is repeated for a predefined number of iterations. Another important aspect of our framework is that the VQ-VAE model and the MAP-Elites algorithm are completely independent. Hence, the user has the freedom to incorporate different variations of the MAP-Elites algorithm as well as other selection, mutation, and crossover operators.
Updating the model: The final component of our framework is the model update, refining its latent space and codebook. The update of the model occurs every $N_{\text{update}}$ iteration and in an online fashion, meaning that the model never resets. Ultimately, the model will converge, and thus, there will be no further changes in the latent space and codebook. To train the model, we exploit the experiences of our agent, by storing the raw data produced by the solutions in the evolution phase. By utilizing these experiences, we can achieve a latent space that encapsulates the underlying dynamics of our agent as well as of the environment. Hence, the behavioral descriptors and the grid of our behavioral space consist only of feasible behaviors to be produced by the agent, something that would have been impossible to achieve by using a naive clustering algorithm.
After each update of the VQ-VAE model, we retain a new behavioral space, which foreshadows that our current grid and BDs are not aligned. Specifically, this means that the same solution before and after the update of the model is not described by the same BD. To correct this, we need to also update our archive and grid of the behavioral space. Consequently, we set the updated codebook as our new grid, and then we reevaluate the solutions we have in the archive, by encoding their raw data and assigning their BDs to the the closest cell as in the evolution phase. The update of the archive serves a double purpose; our solutions are accurately represented by the behavioral space, and both unfit and less diverse solutions are discarded due to the new refined latent space and codebook.
IV-C Behavioral Space Bounding
@@
While this property is advantageous in many contexts, it introduces a potential issue in our framework. Recall that our goal is to learn a behavioral space, represented by a latent space, that allows us to find solutions that are both diverse and high-performing. According to the manifold hypothesis, behavioral descriptors of diverse solutions will naturally be far apart in the latent space. However, the VQ-VAE model, by design, updates its codebook less effectively than the latent space itself. This disparity can result in an archive filled with high-performing solutions that lack diversity. Additionally, training the model in an online manner also introduces a lot of difficulties, since, until the model converges, newly found diverse solutions are actually out-of-distribution, likely leading to their latent representation not being representative.
To address this challenge, we constrain the latent (and therefore the behavioral) space to $[-1,1]$. This bounded space ensures that the latent space remains continuous and expressive, while preventing excessive dispersion that might hinder the codebook’s ability to form a meaningful discretization of the behavioral space. By maintaining this bounded structure, we enable the model to produce a diverse set of solutions while preserving high performance.
With this bounded latent space in place, we can further exploit our model to accelerate convergence. Initially, both the VQ-VAE model and its codebook are randomly initialized, which typically requires a significant number of gradient steps to achieve a well-structured discretization. However, given the fixed bounds of the latent space ($[-1,1]$), we can initialize the codebook more effectively. Specifically, we pre-train the codebook by fitting the K-Means algorithm on a sufficient number of samples drawn from the uniform distribution, $\mathcal{U}_{[-1,1]}$.
This initialization provides the codebook with centroids that are already representative of the bounded latent space. As a result, the model requires fewer gradient updates to achieve an optimal discretization of the behavioral space. This accelerates both the learning process and the creation of a well-structured archive, ensuring that the algorithm efficiently identifies diverse, high-quality solutions.
@@ IV-D Cooperation over Competition
During the evolution phase, new solutions are introduced to the archive by competing with existing ones. However, in the early iterations, the archive often contains less diverse solutions because the representation learning model has not yet fully established a well-defined behavioral space. As a result, early individuals in the archive may achieve high fitness scores but lack diversity.
To address this issue, we introduce a cooperation aspect to our algorithm during the initial iterations. This cooperation neglects the competition of individuals, and instead, the grid is filled with the most recent solution assigned to each cell, regardless of fitness. In a way, the individuals are cooperating with offsprings of the population, with the goal of promoting the diversity of the archive. This approach allows the model to focus on constructing a highly diverse grid or refining the behavioral space before strict competition begins. Additionally, this strategy enhances the algorithm’s exploration capabilities by ensuring frequent updates to the solution set, promoting diversity during the critical early stages of the evolutionary process. We should emphasize that the VQ-Elites is robust and did not require the cooperation component for our experiments, yet we found it critical for other unsupervised QD algorithms we evaluated.
@@ V Experiments
For our experiments, we selected two state-covering tasks using a mobile robot and two pose-reaching tasks with a 7 degrees of freedom robotic arm. These tasks were chosen to showcase the versatility and effectiveness of VQ-Elites in solving robotics problems without requiring prior knowledge of the environment or the system. Moreover, these experiments demonstrate the broader potential of VQ-Elites beyond these specific scenarios.
A key advantage of VQ-Elites is its minimal reliance on hyperparameter tuning. Throughout all experiments, the crossover and mutation rates were kept consistent, and the model was updated every five iterations of the evolutionary process. This consistency highlights the robustness and ease of deployment of VQ-Elites across diverse tasks.
@@ V-A Mobile Robot Experiments
$r=-\sum_{i}^{N}||\boldsymbol{p}_{i}-\boldsymbol{p}_{N}||_{2},$ (8)
where $\boldsymbol{p}_{i}$ is the position of the robot in the $i$-th step of its trajectory. For every candidate solution, the raw BD is an image of the last position of the robot in the environment with dimensions $64\times 64$, which we feed into the encoder network of VQ-VAE to extract our low-dimensional BD. Additionally, these images are stored for the learning phase of our pipeline. We should note, that the reward of the environment is only responsible for the fitness of the solution. The space-covering aspect of the task is fulfilled in the learning phase, and thus, the model should construct an optimal desensitization of our behavioral space to achieve it.
@@
Figure 2: Visualization of the obstacle-free and maze environments for the mobile robot experiments.
@@ V-B Robotic Arm Experiments
Regarding the pose-reaching tasks, the general notion of the pipeline follows the mobile robot experiment, with some variations. The goal of the task is for the end-effector of the robot to reach a predefined location, with as diverse joint configurations as possible. Here, the input of our policy as well as the data we use to extract the BD and train the VQ-VAE model are the joints of the robot (raw BD). The output of the policy is the joint velocities.
@@
In order to demonstrate the flexibility of our algorithm, we first showcase this experiment with the default limits of the joint space, and after, we introduce some artificial constraints to the robot, which simulate a problematic agent or a confined operation space. The goal of this variation is to exhibit the high potential of unsupervised learning in combination with QD algorithms in solving problems in an agnostic manner, where the constraints of the system cannot be identified beforehand. For both of these tasks, we use 3000 iterations and each policy performs 300 environment steps.
V-C Evaluation Metrics
@@
To evaluate the performance of our algorithm, we employ most of the metrics used in AURORA [9]. Specifically, we evaluate our approach in terms of coverage, grid mean fitness, and QD score.
V-C1 Coverage
@@
This metric measures the percentage of the hand-coded behavioral space that is covered. The BD space is divided into a grid of cells, and the coverage is calculated as the proportion of cells containing at least one individual. In the mobile robot experiments, since we operate in a 2D space, we simply discretize the robot’s operating space in cell of dimension $N\times N$. On the other hand, this is not possible in the pose-reaching task, hence, we gather 1M poses using inverse kinematics (IK) and perform a K-Means clustering on these data. The resulting centroids correspond to our hard-coded BD.
V-C2 Grid Mean Fitness
This metric assesses the overall quality of a container by calculating the mean fitness of individuals in a $N\times N$ grid of the hand-coded BD space. For each cell, only the individual with the highest fitness is retained, and their fitnesses are averaged to compute the grid mean fitness. This approach ensures that the metric is independent of the solution distribution within the BD space, enabling fair comparisons across algorithms using learned BD spaces.
V-C3 Quality-Diversity Score
@@
      Experiment       Mobile robot       Robotic arm
      Iterations       $3\times 10^{3}$       $3\times 10^{3}$
      Population size       $128$       $128$
      Mutation rate       $10\%$       $30\%$
      Cross-over rate       $10\%$       $20\%$
      $|\mathcal{C}|$       $2000$       $400$
      BD/Latent size       $2$       $5$
      $\mathcal{E}$       $10$       $10$
      $lr$       $7\times 10^{-4}$       $7\times 10^{-4}$
      Batch size       $64$       $64$
      $N_{\text{update}}$       $5$       $5$
      $N_{\text{cooperation}}$       $1\times 10^{3}$       $1\times 10^{3}$
@@ VI Results
In this section, we showcase the results of our algorithm in the experimental evaluation, starting with the mobile robot experiments, then continuing to robotic arm experiments and finally presenting an ablation study we performed on our algorithm.
@@ VI-A Mobile Robot Results
For the mobile robot experiments, we compare VQ-Elites algorithm with the classic MAP-Elites algorithm, the AURORA approach as well as the AURORA after integrating our two proposed components, behavioral space bounding and cooperation, which from now on we will refer as AURORA. We should mention that a variation of this experiment is also present in the AURORA paper [9], hence it is ideal for comparison purposes, however, we are unable to exactly replicate the authors’ results [^1] We used the exact hyperparameters presented in the original paper and communicated with the authors., as demonstrated in the manuscript.
@@
In the case where the robot operates in an obstacle-free environment, we can observe that all the algorithms, except the vanilla AURORA, perform almost perfectly in terms of coverage, assigning solutions to nearly every grid cell (Fig. 3). In terms of the QD score, we see that the classic MAP-Elites algorithm produces more high-quality and diverse solutions compared to the other approaches. Both of these observations are expected, since we can easily define a behavioral grid manually, and still every cell will represent a feasible behavior. Nonetheless, it is evident that algorithms that assume no prior knowledge about the environment such as ours and other unsupervised QD algorithms infused with our cooperation and behavioral space bounding components, can achieve the performance of MAP-Elites without full knowledge of the task (answers to Question 1).
@@
Figure 3: Results for the mobile robot operating in the obstacle-free environment of four different algorithms, VQ-Elites, AURORA, AURORA, and MAP-Elites with hard-coded behavioral descriptors. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
@@
Transitioning to the L-shaped environment, our approach and AURORA outperform the others by a large margin in terms of coverage, while VQ-Elites exhibits the best QD score of all (Fig. 4). This is one type of task that makes the existence of unsupervised QD algorithms essential, since, if we do not possess full knowledge about the task at hand, and particularly in this scenario, we do not know the environment, we cannot construct a grid such that every cell corresponds to a possible behavior (answers to Question 2). On the other hand, a data-driven unsupervised approach is more than capable of constructing such a grid.
@@
Figure 4: Results for the mobile robot operating in the L-shaped environment of four different algorithms, VQ-Elites, AURORA, AURORA, and MAP-Elites with hard-coded behavioral descriptors. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
@@
Lastly, since the experiment is two-dimensional, we also provide plots, which are hard or infeasible to do or comprehend in high dimensions. These plots depict the grid mean fitness, final solutions/archive of our algorithm, and the behavior descriptors and behavioral grid for each scenario in Fig. 5. In the left plot, we illustrate the grid mean fitness by discretizing the robot’s operating space into a $8\times 8$ grid, and the color of each cell represents the normalized reward/fitness of the best solution assigned to each cell. The middle plot showcases the actual solutions VQ-Elites produced, while the right plot depicts the BDs of the solutions that are shown in the middle plot along with all the centers of our behavioral space grid learned by the VQ-VAE model. Overall, we observe that VQ-Elites is able to effectively explore the behavior space, produce meaningful latent representations, and automatically perform clustering of the elite behaviors.
Figure 5: Plots of the mean fitness in an 8x8 grid, final archive of solutions, and the behavior descriptors and learned behavioral grid of: (a) obstacle-free environment. (b) L-shaped environment.

### v2

@@ IV-B Evolution Phase
The next key component of VQ-Elites is the evolution phase. During this phase the Quality-Diversity algorithm is executed, performing the usual steps. Foremost, the archive is randomly initialized with random solutions. Then, a random batch of solutions is selected, which are alternated through mutation and crossover. The new variation of the solutions are executed, and evaluated through their behavior descriptor. In order to extract these BDs, we encode the raw data produced by our solutions with the encoder of our VQ-VAE model, acquiring the corresponding continuous latent vectors. For each BD, we find the closest grid cell, and if the cell is empty the solution is added to the archive, otherwise, we compare the fitness of the current solution with the one of the solution that is already in the archive. The solution in the archive is replaced only if the new solution is fitter. This procedure is repeated for a predefined number of iterations. Another important aspect of our framework is that the VQ-VAE model and the MAP-Elites algorithm are completely independent. Hence, the user has the freedom to incorporate different variations of the MAP-Elites algorithm as well as other selection, mutation, and crossover operators.
Updating the model: The final component of our framework is the model update, refining its latent space and codebook. The update of the model occurs every $N_{\text{update}}$ iteration and in an online fashion, meaning that the model never resets. Ultimately, the model will converge, and thus, there will be no further changes in the latent space and codebook. To train the model, we exploit the experiences of our agent, by storing the raw data produced by the solutions in the evolution phase. By utilizing these experiences, we can achieve a latent space that encapsulates the underlying dynamics of our agent as well as of the environment. Hence, the behavior descriptors and the grid of our behavior space consist only of feasible behaviors to be produced by the agent, something that would have been impossible to achieve by using a naive clustering algorithm.
After each update of the VQ-VAE model, we retain a new behavior space, which foreshadows that our current grid and BDs are not aligned. Specifically, this means that the same solution before and after the update of the model is not described by the same BD. To correct this, we need to also update our archive and grid of the behavior space. Consequently, we set the updated codebook as our new grid, and then we reevaluate the solutions we have in the archive, by encoding their raw data and assigning their BDs to the closest cell as in the evolution phase. The update of the archive serves a double purpose; our solutions are accurately represented by the behavior space, and both unfit and less diverse solutions are discarded due to the new refined latent space and codebook.
IV-C Behavior Space Bounding
@@
While this property is advantageous in many contexts, it introduces a potential issue in our framework. Recall that our goal is to learn a behavior space, represented by a latent space, that allows us to find solutions that are both diverse and high-performing. According to the manifold hypothesis, behavior descriptors of diverse solutions will naturally be far apart in the latent space. However, the VQ-VAE model, by design, updates its codebook less effectively than the latent space itself. This disparity can result in an archive filled with high-performing solutions that lack diversity. Additionally, training the model in an online manner also introduces a lot of difficulties, since, until the model converges, newly found diverse solutions are actually out-of-distribution, likely leading to their latent representation not being representative.
To address this challenge, we constrain the latent (and therefore the behavior) space to [1,1][-1,1]. This bounded space ensures that the latent space remains continuous and expressive, while preventing excessive dispersion that might hinder the codebook’s ability to form a meaningful discretization of the behavior space. By maintaining this bounded structure, we enable the model to produce a diverse set of solutions while preserving high performance.
With this bounded latent space in place, we can further exploit our model to accelerate convergence. Initially, both the VQ-VAE model and its codebook are randomly initialized, which typically requires a significant number of gradient steps to achieve a well-structured discretization. However, given the fixed bounds of the latent space ([1,1][-1,1]), we can initialize the codebook more effectively. Specifically, we pre-train the codebook by fitting the K-Means algorithm on a sufficient number of samples drawn from the uniform distribution, $\mathcal{U}_{[-1,1]}$.
This initialization provides the codebook with centroids that are already representative of the bounded latent space. As a result, the model requires fewer gradient updates to achieve an optimal discretization of the behavior space. This accelerates both the learning process and the creation of a well-structured archive, ensuring that the algorithm efficiently identifies diverse, high-quality solutions.
@@ IV-D Cooperation over Competition
During the evolution phase, new solutions are introduced to the archive by competing with existing ones. However, in the early iterations, the archive often contains less diverse solutions because the representation learning model has not yet fully established a well-defined behavior space. As a result, early individuals in the archive may achieve high fitness scores but lack diversity.
To address this issue, we introduce a cooperation aspect to our algorithm during the initial iterations. This cooperation neglects the competition of individuals, and instead, the grid is filled with the most recent solution assigned to each cell, regardless of fitness. In a way, the individuals are cooperating with offsprings of the population, with the goal of promoting the diversity of the archive. This approach allows the model to focus on constructing a highly diverse grid or refining the behavior space before strict competition begins. Additionally, this strategy enhances the algorithm’s exploration capabilities by ensuring frequent updates to the solution set, promoting diversity during the critical early stages of the evolutionary process. We should emphasize that the VQ-Elites is robust and did not require the cooperation component for our experiments, yet we found it critical for other unsupervised QD algorithms we evaluated.
@@ V Experiments
  5. 5.
    Can VQ-Elites solve hard exploration tasks?
  6. 6.
    Can VQ-Elites operate with high-dimensional policy structures?
For our experiments, we selected two state-covering tasks using a mobile robot, two pose-reaching tasks with a 7 degrees of freedom robotic arm, and a hard exploration and discrete task inspired by the MiniGrid Reinforcement Learning environment [36]. These tasks were chosen to showcase the versatility and effectiveness of VQ-Elites in solving problems without requiring prior knowledge of the environment or the system. Moreover, these experiments demonstrate the broader potential of VQ-Elites beyond these specific scenarios.
A key advantage of VQ-Elites is its minimal reliance on hyperparameter tuning. In most of the experiments, the crossover and mutation hyperparameters were kept consistent, as well as the model’s update frequency and number of training epochs. This consistency highlights the robustness and ease of deployment of VQ-Elites across diverse tasks.
@@ V-A Mobile Robot Experiments
$f=-\sum_{i}^{N}\|\boldsymbol{p}_{i}-\boldsymbol{p}_{N}\|_{2},$ (8)
where $\boldsymbol{p}_{i}$ is the position of the robot in the $i$-th step of its trajectory. For every candidate solution, the raw BD is an image of the last position of the robot in the environment with initial dimensions $600\times 600$, which is then downscaled to $64\times 64$ and fed into the encoder network of VQ-VAE to extract our low-dimensional BD. Additionally, these images are stored for the learning phase of our pipeline. We should note, that the reward of the environment is only responsible for the fitness of the solution. The space-covering aspect of the task is fulfilled in the learning phase, and thus, the model should construct an optimal desensitization of our behavior space to achieve it.
@@
Figure 2: Visualization of the obstacle-free and maze environments for the mobile robot experiments, as well as the starting positions of the robot.
@@ V-B Robotic Arm Experiments
Regarding the pose-reaching tasks, the general notion of the pipeline follows the mobile robot experiment, with some variations. The goal of the task is for the end-effector of the robot to reach a predefined location, with as diverse joint configurations as possible. Here, the input of our policy as well as the data we use to extract the BD and train the VQ-VAE model are the joints of the robot (raw BD). The output of the policy is the joint velocities. The fitness formula for the robotic arm is:
$f=e^{(-\|E_{g}-E_{N}\|_{2})}$ (9)
where $E_{g}$ is the 3D coordinates of the goal, and $E_{N}$ is the 3D coordinates of our end-effector at the last step of the rollout.
@@
In order to demonstrate the flexibility of our algorithm, we first showcase this experiment with the default limits of the joint space, and after, we introduce some artificial constraints to the robot, which simulate a problematic agent or a confined operation space. The goal of this variation is to exhibit the high potential of unsupervised learning in combination with QD algorithms in solving problems in an agnostic manner, where the constraints of the system cannot be identified beforehand. For both of these tasks, we use 3000 iterations, and each policy performs 300 environment steps.
V-C MiniGrid Exploration
The MiniGrid experiment (Fig. 3) consists of a grid map, and an agent moving (with discrete actions) inside it. We use this experiment to investigate the performance of our approach in a domain other than robotics. The ultimate goal of the task is to discover individuals through the evolution process to traverse all possible tiles of the map. The agent has a field of view (FOV) of 3 tiles and can perform 5 actions, $\mathcal{A}=${up, right, down, left, stay still}, which are selected by the policy. For the policy’s input, we use snapshots of the current agent’s position while respecting its FOV, along with the list of keys that have been gathered using one-hot encoding. Hence, these snapshots contain information in pixels inside the FOV of the agent, while the rest are masked out. Additionally, the snapshot of the last step in the environment also serves as our raw BD. The size of the snapshots is $18\times 18$.
At every episode, the agent needs to solve two subtasks multiple times in order to get access to the rest of the map. Firstly, it needs to identify any keys present and reach their tiles to grab them. Secondly, it needs to find the doors that correspond to those keys to unlock them. By solving these subtasks, the agent is rewarded with 10 points each time it finds a key and with 20 points each time it opens a door. The accumulated points of each rollout act as our fitness.
The two key characteristics that make this experiment interesting are the discrete nature of the environment and its hard-to-explore nature. Compared to the previous experiments, where they are governed by smooth dynamics and transitions, the discrete nature of the current experiment showcases that our algorithm is agnostic to the dynamics of the task at hand. Additionally, this task is a hard-to-explore problem, both in the state space and solution space. To make the state space exploration difficult, we designed a map where the agent needs to solve multiple rooms to traverse the whole map (Fig. 3). On the other hand, the policy we deploy is a CNN network with a high number of parameters ($\approx 3$K), which makes the evolution more hectic and the discovery of optimal solutions harder. We should note that on Fig 3, the graphics are only for visual presentation. In the map used for our experiment, the key, doors, obstacles, and agent are represented by uniquely colored tiles for each item. Overall, for this experiment, we perform 10,000 iterations, and 200 environment steps per evaluation.
Figure 3: Overview of the MiniGrid task’s map. The color-coded circles around the keys and the doors, represent key-door pairings.
V-D Evaluation Metrics
@@
To evaluate the performance of our algorithm, we employ most of the metrics used in AURORA [9], but we also introduce two new metrics. Specifically, we evaluate our approach in terms of coverage, grid mean fitness, QD score, and our newly proposed Effective Diversity Ratio and Coverage Diversity Score.
V-D1 Coverage
@@
This metric measures the percentage of the hand-coded behavior space that is covered. The BD space is divided into a grid of cells, and the coverage is calculated as the proportion of cells containing at least one individual. In the mobile robot experiments, since we operate in a 2D space, we simply discretize the robot’s operating space in cell of dimension $N\times N$. Similarly, in the MiniGrid case, since the state is discrete, we can easily measure the ground truth coverage. On the other hand, this is not possible in the pose-reaching task, hence, we gather 1M poses using inverse kinematics (IK) and perform a K-Means clustering on these data. The resulting centroids correspond to our hard-coded BD.
V-D2 Quality-Diversity Score
@@
V-D3 Grid Mean Fitness
This metric assesses the overall quality of a container by calculating the mean fitness of individuals in a grid of the hand-coded BD space. For each cell, only the individual with the highest fitness is retained, and their fitnesses are averaged to compute the grid mean fitness. This approach ensures that the metric is independent of the solution distribution within the BD space, enabling fair comparisons across algorithms using learned BD spaces. This metric is a normalized version of the QD score, and is basically a qualitative visual metric. As such we apply it only in the mobile robot and MiniGrid scenarios.
V-D4 Effective Diversity Ratio
We introduce the Effective Diversity Ratio (EDR), a novel metric designed to enable fairer comparisons between unsupervised QD algorithms. EDR quantifies the proportion of genuinely diverse solutions discovered by an algorithm, by measuring the ratio of unique behavior outcomes to the total number of valid individuals in the archive. To compute EDR, we discretize the behavior space into $N$ bins, where $N$ represents the number of desired distinct behaviors. Each individual in the archive is evaluated by executing it in the environment (or task), and the resulting solution, i.e., the observed behavior outcome, is assigned to the nearest bin in the discretized space. The EDR is then defined as the number of occupied bins in the ground truth grid divided by the total number of valid individuals in the original archive:
$\text{EDR}=\frac{\text{Number of Occupied Bins}}{\text{Valid Archive Size}}$
This metric reflects how effectively the algorithm explores the behavior outcome space, discounting redundancy among individuals whose executions lead to similar behaviors. A higher EDR indicates broader behavior coverage and less repetition in the discovered solutions.
V-D5 Coverage Diversity Score
We also introduce the Coverage Diversity Score (CDS), a metric designed to assess behavioral diversity in the context of archive coverage. It is defined as the product of Coverage and EDR:
$\text{CDS}=\text{Coverage}\;\times\;\text{EDR}$
While EDR captures the raw diversity of the archive, it can be misleading in edge cases. For example, at the very beginning of evolution, an archive containing a single individual will yield an EDR of $100\%$, despite offering no meaningful diversity. Conversely, as the archive grows, the EDR may become artificially low even if the behavior spread remains satisfactory. By incorporating coverage, the CDS mitigates these issues, providing a more balanced view of diversity relative to the size of the explored behavior space. It offers a more robust indicator of how effectively an algorithm populates the behavior space with both broad and distinct solutions.
@@ V-D5 Coverage Diversity Score
Experiment Mobile robot Robotic arm MiniGrid
Iterations $3\times 10^{3}$ $3\times 10^{3}$ $10^{4}$
Population size $128$ $128$ $128$
$|\mathcal{C}|$ $2000$ $1500$ $400$
Raw BD size $[1,64,64]$ 66 $[3,18,18]$
BD/Latent size 22 55 55
$\mathcal{E}$ $10$ $10$ $10$
lrlr $7\times 10^{-4}$ $7\times 10^{-4}$ $7\times 10^{-4}$
Batch size $64$ $64$ $64$
$N_{\text{update}}$ 55 55 $10$
$N_{\text{cooperation}}$
@@ VI Results
In this section, we showcase the results of our algorithm in the experimental evaluation, starting with the mobile robot experiments, then continuing to robotic arm experiments and the MiniGrid exploration task, and finally presenting ablation studies we performed on our algorithm.
@@ VI-A Mobile Robot Results
For the mobile robot experiments, we compare VQ-Elites algorithm with the classic MAP-Elites algorithm, the AURORA approach, as well as the AURORA after integrating our two proposed components, behavior space bounding and cooperation, which from now on we will refer to as AURORA. We should mention that a variation of this experiment is also present in the AURORA paper [9], hence it is ideal for comparison purposes, however, we are unable to exactly replicate the authors’ results [^2] We used the exact hyperparameters presented in the original paper and communicated with the authors., as demonstrated in the manuscript.
@@
In the case where the robot operates in an obstacle-free environment, we can observe that all the algorithms, except the vanilla AURORA, perform almost perfectly in terms of coverage, assigning solutions to nearly every grid cell (Fig. 4). In terms of the QD score, we see that the classic MAP-Elites algorithm produces more high-quality and diverse solutions compared to the other approaches. Both of these observations are expected, since we can easily define a behavior grid manually, and still every cell will represent a feasible behavior. Nonetheless, it is evident that algorithms that assume no prior knowledge about the environment, such as ours and other unsupervised QD algorithms infused with our cooperation and behavior space bounding components, can achieve the performance of MAP-Elites without full knowledge of the task (answers to Question 1). Additionally, considering the EDR and CDS metrics, the VQ-Elites is producing slightly more diverse behaviors compared to MAP-Elites and AURORA. This indicates that despite the coverage of VQ-Elites being lower compared to the other two, the behaviors produced do a better job of promoting diversity while maintaining a high-quality archive.
@@
Figure 4: Results for the mobile robot operating in the obstacle-free environment of four different algorithms, VQ-Elites, AURORA, AURORA, and MAP-Elites with hard-coded behavior descriptors. Each graph line represents the median and the 25th and 75th quantile over ten runs for the corresponding algorithm.
@@
Transitioning to the L-shaped environment, our approach and AURORA outperform the others by a large margin in terms of coverage, while VQ-Elites exhibits the best QD score of all (Fig. 5). This is one type of task that makes the existence of unsupervised QD algorithms essential, since, if we do not possess full knowledge about the task at hand, and particularly in this scenario, we do not know the environment, we cannot construct a grid such that every cell corresponds to a possible behavior (answers to Question 2). On the other hand, a data-driven unsupervised approach is more than capable of constructing such a grid. Although MAP-Elites has low coverage, it outperforms the other algorithms in the effective diversity metric. This is expected since the hand-crafted behavior grid is well-discretized, enabling the algorithm to provide a small but diverse archive. However, this is not the case in the CDS metric, where VQ-Elites has achieved a higher score since the coverage is taken into account. Additionally, between VQ-Elites and AURORA, our algorithm performs better in the EDR and CDS metrics, since it does not rely on an unstructured archive, which populates the archive based on heuristics that may result in many duplicate behaviors.
@@
Figure 5: Results for the mobile robot operating in the L-shaped environment of four different algorithms, VQ-Elites, AURORA, AURORA, and MAP-Elites with hard-coded behavior descriptors. Each graph line represents the median and the 25th and 75th quantiles over ten runs for the corresponding algorithm.
@@
Lastly, since the experiment is two-dimensional, we also provide plots, which are hard or infeasible to do or comprehend in high dimensions. These plots depict the grid mean fitness, final archives of the algorithms, and the behavior descriptors for each algorithm per scenario in Fig. 6. In the first column of each plot, we illustrate the grid mean fitness by discretizing the robot’s operating space into an $8\times 8$ grid, and the color of each cell represents the normalized fitness of the best solution assigned to each cell. The second column showcases the actual solutions VQ-Elites produced, while the last one depicts the BDs of the solutions that emerged. Overall, we observe that VQ-Elites and AURORA are able to effectively explore the behavior space, produce meaningful latent representations, and in the case of VQ-Elites, it automatically performs clustering of the behaviors, constructing a meaningful behavior space grid.
Figure 6: Visualizations that demonstrate the mobile experiment’s outcomes of the three algorithms in the two different environments. Left: Obstacle-free experiment outcome. Right: L-Shaped experiment outcome. In both plots, the first column indicates the grid mean fitness in a $8\times 8$ grid, while the second depicts where every learned behavior ends up in the environment. The third column is a visualization of the behavior descriptors that describe these behaviors.